2. 1500 famil. ies wuh 2 children were sekcted randomly, and the following data were recorded: •

Number of girls in a family 2 1 0
Num_ber of families 475 814 211


Compute the probability of a family, • chosen at random,
having


www .cbsepdf .com


232 1\11ATHEMATICS-IX
(i) 2 girls (ii) 1 girl (iii ) No girL Also check whether the sum of these probabilities is 1.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Tot.al number of families = 475 + 814 + 211 = 1500. (i) Probability or a fmDilies having 2 girls. = Number of families havi-n-g'-2--g-irls number of families 475 19 = - = - 1600 60 (ii) Probability of a families having 1girl = Number of families having 1girl Total number of families 814 407 = 1 600 = 750 (ii.i) Probability of family having no girl Number of families havingno girl 211 = Tot.al nwnber of families• - 1 500 • . . . 476 814 211 Sum of the probability = 1500 +. 1500 + 1500 476 + 814 +211 1600 = 1600 = 1 600 = 1 Clearly, the sum of the three probabilities is 1.